### v1

Abstract
Multimodal data has become a crucial element in the realm of big data analytics, driving advancements in data exploration, data mining, and empowering artificial intelligence applications. To support high-quality retrieval for these cutting-edge applications, a robust data retrieval platform should meet the requirements for transparent data storage, rich hybrid queries, effective feature representation, and high query efficiency. However, among the existing platforms, traditional schema-on-write systems, multi-model databases, vector databases, and data lakes, which are the primary options for multimodal data retrieval, are difficult to fulfill these requirements simultaneously. Therefore, there is an urgent need to develop a more versatile multimodal data retrieval platform to address these issues.
In this paper, we introduce a Multimodal Data Retrieval Platform with Query-aware Feature Representation and Learned Index based on Data Lake (MQRLD). It leverages the transparent storage capabilities of data lakes, integrates the multimodal open API to provide a unified interface that supports rich hybrid queries, introduces a query-aware multimodal data feature representation strategy to obtain effective features, and offers high-dimensional learned indexes to optimize data query. We conduct a comparative analysis of the query performance of MQRLD against other methods for rich hybrid queries. Our results underscore the superior efficiency of MQRLD in handling multimodal data retrieval tasks, demonstrating its potential to significantly improve retrieval performance in complex environments. We also clarify some potential concerns in the discussion.
keywords:
[…]
{highlights}
The study offers a multimodal data retrieval platform, which supports transparent data storage by using data lake technology, enables rich hybrid queries through the usage of a multimodal open API, and provides a query-aware mechanism to optimize retrieval.
The study proposes a multimodal feature representation technique that transforms raw multimodal data into representative features, improving effective retrieval of multimodal data.
The study introduces a high-dimensional learned index that can adaptively optimize its inner structure, enhancing efficient retrieval of multimodal data.
1 Introduction
Multimodal data (including video, audio, text, etc.) has reached a critical mass, with 80 percent of worldwide data expected to be in these modalities by 2025 (King, 2020). This variety and sheer volume make "Multimodal Data Retrieval" a core discipline of the more general domain of "Big data" research, which is the cornerstone for supporting big data analysis, driving advancements in data exploration, data mining, and empowering artificial intelligence applications. The efficient handling of large-scale multimodal data is crucial for digging deeper insights and driving innovation across diverse domains, including multimodal knowledge graph (Zhang et al., 2020; Abu-Salih, 2021), personalized recommendation (Wu et al., 2024), disease prediction in healthcare (Andorra et al., 2024; Thukral et al., 2023), and fake news detection in law enforcement (Xue et al., 2021), etc. Moreover, with the evolution and advancement of large-scale models, especially multimodal large-scale language models, AI is endowed with the ability to accept multi-sensory inputs like humans and provide more human-like interactions, which relies heavily on the multimodal data retrieval platform.
The main challenge of a multimodal data retrieval platform is how to efficiently and accurately exploit valuable information from the vast and diverse multimodal data through rich query options along with the concern of storage and management of fast-growing data (Mishra and Misra, 2017). For instance, as shown in Fig 1, we possess multimodal data that includes various types of information such as text, images, video, audio, etc. This data might encompass sports news, dance videos, and life recordings. Our objective is to quickly retrieve the desired information from vast and diverse multimodal data through querying. This requires a retrieval platform that combines efficient storage and querying capability, meeting the following requirements:
Figure 1: Querying in Multimodal Data
(1)Storage layer: Transparent data storage. Multimodal data is typically unstructured and lacks a fixed organization or predefined format, which makes it challenging to store and access data in a seamless manner without requiring users to understand the underlying complexities (Gao et al., 2020). For instance, multimodal data consists of data stored in various formats such as JPG, TXT, MKV, and MP3 files, making the management of such data considerably difficult to hide from the end user.
(2)Query layer: Rich hybrid queries and query-aware mechanism. The effectiveness of information mining for multimodal data largely depends on the richness of the query capabilities provided. Hybrid queries (Wei et al., 2020) refers to search operations over a single vectorized attribute and a set of numerical attributes associated with the searching object, whereas rich hybrid queries extend this concept by supporting more flexible combinations of a broader spectrum of "attribute", which can be vectors or numeric values (see details in Section 4.2). For instance, in the query "Find the information from the soccer match on August 10, 2014, including the images and audio. During the match, the African players in red jerseys repeatedly dribbled through the Asian players in white, earning loud cheers from the audience.", "August 10, 2014" serves as a numeric query, while the subsequent sentence functions as a vector query. Moreover, query-aware refers to the logs of the rich hybrid queries, which involve both execution processes and query results. By maintaining these records, systems can gain insights into query behaviors, which can be used to optimize the effectiveness and efficiency of future multimodal data retrieval operations.
(3) Feature representation layer: Effective feature representation. A common approach to multimodal data retrieval is to "understand" multimodal data by representing it as features and using semantic search to obtain data with corresponding features (Wang et al., 2016). Yet, different multimodal datasets possess diverse and distinct characteristics, and varying query contexts may emphasize different features. Therefore, it is challenging to employ a static process to transform raw data into optimal features and data layouts to meet the requirements of rich hybrid queries. Instead, a query-aware process is necessary to obtain the optimal features.
(4)Index layer: High query efficiency. Different users and application scenarios usually have diverse query patterns. Journalists, wish to quickly retrieve sports news from the multimodal data, whereas for bloggers, their queries are more likely related to life recordings. It has emerged as a significant challenge to devise and update system’s index structure to adapt to different query workloads and efficiently manage computational costs.
In response to the growing need for inferring meaningful insights from large amounts of multimodal data, data platforms have been evolving to meet the demands of multimodal data retrieval. Existing retrieval platforms can be categorized into traditional schema-on-write systems, multi-model databases, vector databases, and data lakes. (1) Traditional schema-on-write (i.e., store data in predefined schema) systems, like relational databases, are inflexible for handling multimodal data transparent storage and hybrid queries due to their fixed schema constraints when managing data. (2) Multi-model databases have been proposed to integrate several schema models within a single, integrated backend, allowing them to handle several data formats within one unified system (Lu and Holubová, 2019; Lei et al., 2024). While these databases can offer transparent storage for certain data modalities (e.g., text), their storage capabilities are often limited when dealing with a comprehensive range of multimodal data. Additionally, their functionality and efficiency are constrained by the lack of rich query options and robust indexing structure. (3) Vector databases are designed specifically for efficient retrieval of high-dimensional vector data through vector similarity searches or simple hybrid queries (Pan et al., 2024), often overlooking the storage of the original multimodal data and the process of feature representation. Despite their specialized design, vector databases face significant challenges in maintaining efficiency when managing vectors with higher dimensionality (Taipalus, 2024). (4) Data lakes are repositories with data from disparate sources that are transparently stored in their original, raw format (Khine and Wang, 2018; Hai et al., 2023). Organizations often regard data lakes as a cost-effective solution for storing vast amounts of data and enabling additional computational functionalities to support retrieval operations, but most of them lack support for rich hybrid queries (Ren et al., 2021b, a; Xiao et al., 2022). Furthermore, the efficiency of queries in data lakes is often compromised due to the insufficient focus on the feature representation and indexing processes of raw data. Therefore, in this study, we explore how to leverage the potential of the data lake to propose a more full-fledged multimodal data retrieval platform based on data lake. In addition to providing infrastructural transparent storage and rich hybrid queries, we mainly focus on improving the effectiveness and efficiency of the data retrieval platform.
For improving effectiveness, the robustness of the feature representation is critical. This involves transforming raw data into representative features and optimal data layout, in order to benefit subsequent tasks such as indexing and querying. Various embedding approaches have been proposed, ranging from single-modality embedding methods like doc2vec (Le and Mikolov, 2014) and graph2vec (Grohe, 2020) to advanced multi-modality embedding techniques such as CLIP (Radford et al., 2021). Despite the availability of numerous embedding methods, it remains a challenging task to measure and enhance these features to make them semantically rich and computationally efficient for retrieval tasks. Existing feature measurement and enhancement methods primarily focus on optimizing features for building machine learning models. Feature measurement techniques aim to identify features best suited for model construction. Feature enhancement methods, such as imputation (Awawdeh et al., 2022), discretization (Dougherty et al., 1995), and feature creation (Rohrer, 2011; Falcon et al., 2022), target at improving model training performance. These methods are not tailored to the scenario of multimodal data retrieval where the goal is enabling the multimodal retrieval platform to perform queries with greater accuracy and speed.
For improving efficiency, the quality of indexes is crucial for achieving accurate and fast data retrieval. (1) Traditional multi-dimensional indexes such as grid file (Nievergelt et al., 1984) and R-tree (Guttman, 1984) are designed to efficiently support numeric queries, particularly on low-dimensional data, while having difficulties when faced with high-dimensional data. (2) Vector similarity indexes such as ANNOY (Bernhardsson, 2015) and HNSW (Malkov and Yashunin, 2018), excel in the realm of high-dimensional vector data indexing through distance computation in vector space to find similar vectors. Nevertheless, they cannot handle hybrid queries that combine numeric queries and vector queries, and increasing data dimensionality and volume can significantly reduce their query efficiency. (3) High-dimensional learned indexes extends traditional multi-dimensional structures by employing machine learning models to quickly and accurately find expected results (Al-Mamun et al., 2024). This method offers a pathway for indexes to optimize their internal structures by learning data characteristics and query behaviors, especially in scenarios involving large datasets. But still, current works in learned indexes primarily focus on proving efficiently low-dimensional numeric queries. So far, exploring an efficient query method that supports both vector and numeric queries, as well as their combinations, remains a challenge yet to be fully resolved.
Therefore, we summarize the key problems to be solved in this paper as:
  • 1.
    How to achieve transparent and seamless storage, as well as rich hybrid queries, for multimodal data retrieval?
  • 2.
    How to convert multimodal data into an optimal feature representation to enhance query result quality?
  • 3.
    How to support efficient multimodal data retrieval?
To address these problems, we present a multimodal data retrieval platform with query-aware feature representation and learned index based on data lake (MQRLD), which supports transparent data storage, and rich hybrid queries empowered by our high-dimensional learned index that benefits from effective multimodal data feature representation and a query-aware mechanism to deliver optimal performance. The main contributions are as follows:
  1. 1.
    We provide a seamless management and flexible access method for multimodal data, including the utilization of data lake technology to support multimodal data storage, and the usage of multimodal open API to support rich hybrid queries, based on which an efficient query-aware mechanism is built to enhance query performance.
  2. 2.
    We propose a multimodal data feature representation technique designed to transform multimodal raw data into representative features and optimal data layout by perceiving both data characteristics and query behaviors, in order to facilitate index building, thereby enhancing the query effectiveness.
  3. 3.
    We present a high-dimensional learned index that can adaptively optimize its inner structure under different feature data layouts and various task scenarios to support flexible and efficient data retrieval of multimodal data.
2 Related Work
In view of the requirements mentioned above, we summarize the current works on data retrieval platforms, feature representation techniques, and indexes. We first enumerate platforms supporting multimodal data retrieval in Section 2.1, focusing on architecture and functionality which include storage, query options, feature representation functionality, and index type. Then in Section 2.2, we introduce the existing techniques in feature representation domain. Finally, we conduct a more detailed comparison of high-dimensional indexes in Section 2.3, focusing on functionality and performance, which include index structure, query types, and whether data-aware (considering multimodal data characteristics) or query-aware (considering query behaviors) techniques are applied to improve index efficiency.
2.1 Related work on data retrieval platform
Table 1 lists existing data retrieval platforms that support multimodal data retrieval. We can classify them into traditional schema-on-write systems, multi-model databases, vector databases, and data lakes.
Table 1: Data Retrieval Platforms.
Storage Layer Query Layer Feature Representation Layer Index Layer
Name Type Transparent Data Storage Rich Hybrid Queries Queryaware Mechanism Embedding Measurement Enhancement Index
Ours data lake high-dimensional learned index
PostgreSQL (Obe and Hsu, 2017) schema-on-write system traditional multi-dimensional index/ vector similarity index
ArangoDB
(Belgundi et al., 2023) multi-model database traditional multi-dimensional index
Azure Cosmos DB (Guay Paz, 2018) multi-model database traditional multi-dimensional index
OrientDB (Ritter et al., 2021) multi-model database traditional multi-dimensional index
Faiss (Johnson et al., 2019) vector database vector similarity index
Pinecone (Pinecone, 2024) vector database vector similarity index
ADAMpro
(Giangreco and Schuldt, 2016) vector database vector similarity index
Cottontail DB (Gasser et al., 2020) vector database vector similarity index
Milvus (Wang et al., 2021) vector database vector similarity index
HMDFF (Ren et al., 2021b) data lake traditional multi-dimensional index
MHDP (Ren et al., 2021a) data lake N/A
MHDML (Xiao et al., 2022) data lake N/A
Traditional schema-on-write systems require a fixed schema for data storage, organize numerical data in columns, and utilize traditional one-dimensional or multi-dimensional indexes for data retrieval. This structure is well-suited for managing numerical data, for example, PostgreSQL performs effectively in supporting bank transactions (Roskladka et al., 2019). However, this fixed schema approach is less adaptable to high-dimensional data. Although plugins like pgvector (pgvector, 2024) have been introduced to handle vector-level data processing and similarity search, they are unable to deliver high performance and scalability required for large-scale applications.
Multi-model databases aim to accommodate different data types by designing separate schema models (e.g., graph model for network data, document model for JSON and XML data, key-value model for ID-identified data, etc.). While this targeted design provides centralized management for various data types, it lacks generality and struggles to accommodate the increasing variety of data modalities. Furthermore, although most multi-model databases can support simple hybrid queries such as text content search combined with numeric search, they are typically confined to specific data modalities. Examples include ArangoDB (Belgundi et al., 2023), OrientDB (Ritter et al., 2021), and CosmosDB (Guay Paz, 2018). Due to the absence of feature representation functionality, these databases are incapable of supporting semantic searches across different data modalities, thus limiting their effectiveness in multimodal data retrieval.
Standard vector databases, such as Faiss (Johnson et al., 2019) and Pinecone (Pinecone, 2024), focus solely on the storage and retrieval of vector-format data and use only vector similarity indexes. This limitation prevents them from executing hybrid queries. Extensions like ADAMpro (Giangreco and Schuldt, 2016), Cottontail DB (Gasser et al., 2020), and Milvus (Wang et al., 2021) offer additional functionalities, such as hybrid queries and feature embedding. However, they exhibit significant flaws: Firstly, they lack explicit processing from raw data to vectors and neglect the storage of raw data. Secondly, despite providing basic feature embedding functionality, they cannot guarantee the quality of features, thus impacting search effectiveness. Thirdly, although they can provide simple hybrid queries, this comes at the expense of significantly reduced query performance. This heavy query process leads to exponential degradation in performance as the vector dimension increases, making it difficult to handle more complex queries.
Data lake (J., 2010), introduced around 2010, is more flexible for different types of analyses, as no decisions regarding the modeling and processing of data have to be made in advance. Due to their low operational costs, high scalability, and flexibility, distributed file systems or object storages, such as the Hadoop Distributed File System (HDFS) (Borthakur et al., 2008), are commonly employed within data lakes for storing raw data (Schneider et al., 2024). However, this flexibility comes at the cost of lower robustness, as the raw data can barely be validated on ingestion. In addition, most data lakes lack structural management of raw data and effective index structure, limiting search performance for large data volumes. Furthermore, many works built on data lake are often tailored for domain-specific data, such as HMDFF (Ren et al., 2021b), MHDP (Ren et al., 2021a), and MHDML (Xiao et al., 2022) focus on medical or healthcare data, lacking generality.
Existing retrieval platforms perform well within their specific application areas; however, none of them are likely to solve all the challenges related to multimodal data retrieval. Moreover, none of these platforms improve query performance using query-aware or learned index techniques, resulting in suboptimal query efficiency.
2.2 Related work on feature representation
Feature representation encompasses the processes of embedding, measurement, and enhancement. Feature embedding, a technique widely studied, has evolved from simple methods like BoW and TF-IDF to advanced techniques such as Word2Vec (Mikolov et al., 2013), GloVe (Pennington et al., 2014), and CLIP (Radford et al., 2021), providing diverse options across different data modalities. These developments have enabled practical applications like YouTube’s content recommendation (Covington et al., 2016). Feature measurement metrics are essential tools to assess the quality, importance, and relevance of embedding features in predictive models. While commonly used metrics like Matthews Correlation Coefficient (MCC), F-value, and variance threshold evaluate a feature’s ability to distinguish between classes, metrics like accuracy and recall reflect overall predictive model performance. However, these metrics often overlook the fidelity of features and fail to account for the specific requirements of query tasks. Various methods have been developed to enhance features and improve model performance, including principal component analysis (PCA), dimensionality reduction techniques (e.g., t-SNE(Cieslak et al., 2020), LDA), scaling, imputation, discretization, and feature creation. Nevertheless, these methods are not specifically tailored to the multimodal data retrieval scenario, where the goal is to perform queries with greater accuracy and speed.
While feature embedding techniques are prevalent in existing data retrieval platforms, few utilize feature measurement and enhancement techniques to optimize downstream query performance. Platforms like ADAMpro and Cottontail DB use numerous feature extractors (Gasser et al., 2020) to generate high-dimensional feature vectors from multimodal data; Milvus leverages deep learning models, including VGG (Simonyan, 2014) and ResNet (He et al., 2016), to convert image data into vectors; HMDFF proposes a multi-aspect embedding method to generate feature vectors representing medical data. Although these platforms support a variety of embedding methods, their selection is often manual and may not be optimal for specific datasets or query workloads. Currently, few of these data retrieval platforms measure and enhance the embedding features to improve their fidelity and generalization, highlighting a gap in ensuring the robustness and adaptability of feature embedding methods for data retrieval scenarios.
2.3 Related work on high-dimensional index
Table 2 lists existing works on high-dimensional indexes. We can categorize them into vector similarity indexes, traditional high-dimensional indexes, and high-dimensional learned indexes.
Vector similarity indexes treat data as points in hyperspace and return the top-k nearest neighbors by comparing "distance" between points. They are primarily classified into three categories: table-based index (e.g., E2 LSH (Datar et al., 2004), IVFADC (Jegou et al., 2010)), tree-based index (e.g., FLANN (Muja and Lowe, 2009), RPTree (Dasgupta and Sinha, 2013)) and graph-based index (e.g., KNNGraph (Dong et al., 2011), HNSW). Although these indexes are effective for vector queries, they struggle to search for single values within the vector, due to their reliance on distance comparison, thus limiting their effectiveness in complex query tasks. Moreover, their query performance declines sharply with an increase in data dimensionality or data records.
Traditional multi-dimensional indexes, such as tree structures (R-tree), grid structures (Grid File), and SFC (space-filling curve)-based structures (z-order (Morton, 1966)), are designed for numeric queries on multi-dimensions, for example, looking up a location on a map by its X and Y coordinates. Their effectiveness is confined to numeric queries on small datasets, as they are functionally unable to support complex queries and their performance degrades significantly on large datasets or high dimensions.
High-dimensional learned indexes have emerged as a prominent topic in recent years. (Kraska et al., 2018) proposed that an index can be treated as a model that learns to locate the expected result, making it particularly suitable for accelerating retrieval in large datasets. Specifically, high-dimensional learned indexes build upon traditional multi-dimensional index structures and construct a learning model to enhance or replace the inner structure of the index by learning data layout (e.g., LISA (Li et al., 2020), RSMI (Qi et al., 2020)) and query workload patterns (e.g., Qd-tree (Yang et al., 2020), Flood (Nathan et al., 2020), Tsunami (Ding et al., 2020)). Although some high-dimensional learned indexes support numeric and vector queries separately (e.g., ML (Davitkova et al., 2020), LISA), they do not support the combination of multiple query types or more complex hybrid queries. Moreover, while these indexes have demonstrated good performance in low-dimensional data, they still suffer from the "curse of dimensionality" phenomenon when applied to higher-dimensional data.
Consequently, the existing high-dimensional indexes still have shortcomings in supporting efficient multimodal data retrieval, including: (1)Lack of Versatile functionality. Existing indexes fail to support rich hybrid query processing such as a combination of numeric and vector queries or multi-vector queries. Although some high-dimensional learned indexes accommodate multiple query types, they execute these queries sequentially, rather than simultaneously. This limitation prevents them from effectively addressing the requirements of complex multimodal queries. (2) Inefficient performance. From Table 1, we can see that most data retrieval platforms adopt vector similarity indexes, and traditional high-dimensional indexes, which lack query-aware capabilities, making them unable to maintain stable performance across different datasets and query workloads. While some platforms provide multiple index structures, they often rely on manual monitoring to select appropriate indexes based on varying query demands. This reliance on human intervention can be impractical and error-prone, particularly in dynamic or large-scale environments where query workloads frequently change. In contrast, learned indexes can adjust their structure based on data-aware mechanisms and query-aware mechanisms. However, they still suffer from the curse of dimensionality when dealing with higher-dimensional data, rendering them unsuitable for multimodal data retrieval scenarios.
Table 2: Multi-dimensional Index.
Structure Learned Data-aware Query-aware Query Type
Ours Tree rich hybrid queries
E2 LSH (Datar et al., 2004) Table vector query
IVFADC (Jegou et al., 2010) Table vector query
FLANN (Muja and Lowe, 2009) Tree vector query
RPTree (Dasgupta and Sinha, 2013) Tree vector query
KGraph (Dong et al., 2011) Graph vector query
HNSW (Malkov and Yashunin, 2018) Graph vector query
ZM (Wang et al., 2019) SFC numeric query
ML (Davitkova et al., 2020) Grid numeric query or vector query
LISA (Li et al., 2020) Grid numeric query or vector query
RSMI (Qi et al., 2020) Grid, SFC numeric query
Qd-tree (Yang et al., 2020) Tree numeric query
Flood (Nathan et al., 2020) Grid numeric query
Tsunami (Ding et al., 2020) Grid, Tree numeric query
COAX (Ghaffari et al., 2020) Grid, Tree numeric query
SPRIG (Zhang et al., 2021) Grid numeric query
PAW (Li et al., 2022) Grid numeric query
LIMS (Tian et al., 2022) Cluster vector query
LMSFC (Gao et al., 2023) SFC numeric query
ELSI (Liu et al., 2023) Tree numeric query
3 Framework
The MQRLD platform efficiently manages and retrieves multimodal data by integrating transparent data storage and supporting rich hybrid queries. It leverages our innovative multimodal data feature representation technique to enhance the learned index structure, thereby improving query performance. As illustrated in Fig 2, the overall framework comprises three core elements: the backbone architecture (Section 4), multimodal data feature representation (Section 5), and high-dimensional learned index (Section 6.1). The backbone architecture consists of three components: a storage foundation for transparent data storage (Section 4.1), a multimodal open API (MOAPI) for rich hybrid queries (Section 4.2), and a query-aware mechanism for further optimization (Section 4.3). This backbone architecture not only provides a standardized storage format and query interface for multimodal data but also preserves quantified records for multimodal query behavior. To further improve retrieval operations, we implement two key methods to flesh out the backbone architecture. First, we employ a multimodal data feature representation technique to transform multimodal data into an optimal format suitable for indexing (Section 5). Second, we design a high-dimensional learned index structure to efficiently retrieve these data (Section 6).
Figure 2: MQRLD Framework
Based on the framework, we introduce the main highlight of each section and the overall workflow of MQRLD, as shown in Fig 3. We use a data lake to seamlessly store and manage multimodal data, serving as the backend storage of our platform (Section 4.1). Building upon it, we provide MOAPI to offer a unified query interface for rich hybrid queries (Section 4.2). While executing each query, we record a series of meaningful statistic data from query behaviors to build the query-aware mechanism (Section 4.3), which is used to refine subsequent feature representation (Section 5) and high-dimensional learned index construction (Section 6). The multimodal data feature representation (Section 5) acts as a bridge between multimodal raw data and the indexes, facilitating the conversion of multimodal raw data into feature vectors that are conducive for index construction through steps of feature embedding, measurement, and enhancement. The embedding and measurement (Section 5.1) aim at discerning representative features for multimodal data on different scenarios. The embedded features then undergo the enhancement process (Section 5.2), which can be seen as an optimization process including transformation and movement in hyperspace based on data patterns and query behaviors, resulting in optimal features and data layout. Finally, the high-dimensional learned index construction (Section 6) fully utilizes the feature representation results to build a cluster tree structure (Section 6.1). To further enhance performance, we optimize index’s inner structure based on query behaviors (Section 6.2), making the index efficient across different query workloads for rich hybrid queries.
Figure 3: MQRLD Workflow
4 MQRLD Backbone Architecture
The MQRLD backbone architecture, based on a data lake, functionally supports multimodal data transparent storage, rich hybrid queries, and query behavior recording. We introduce details of this architecture in the following order: Transparent Storage for Multimodal Data in Section 4.1, MOAPI for Rich Hybrid Queries in Section 4.2, and Statistic Table for Query-aware Mechanism in Section 4.3.
4.1 Transparent Storage for Multimodal Data
Multimodal data retrieval requires storing large amounts of data, each with different modalities that need to be identified and retrieved. Therefore, the primary challenge of storage lies in how to save, manage, and transform this large-scale multimodal data through a transparent process, minimizing the need for manual effort to handle the complexity across various data modalities.
A data lake is a prominent and well-established kind of data platform that stores data in its raw form and object form, thereby preserving all flexible options for future multimodal data retrieval. MQRLD integrates Apache Hudi (Hudi, 2021) as the storage foundation because it is a popular open-source data lake that has been used in many industry applications. One significant advantage of Apache Hudi is its integration with Apache Spark (Shaikh et al., 2019), enabling parallel processing of large-scale data. It utilizes DataFrame, which is conceptually equivalent to a table in a relational database, corresponding to a multimodal object set. As shown in Fig 4, DataFrame columns represent the attributes (such as image texture, text semantics…) of the multimodal data object, embedded by different feature vectors. Each multimodal object can be represented as various feature vectors according to different query tasks, allowing MQRLD to facilitate versatile query tasks across different modalities. For each attribute, the DataFrame also records the embedding models for feature vectors and the HDFS path of the original raw data, enabling quick tracing back to the original data.
Figure 4: Transparent Data Storage and Rich Hybrid Queries in MQRLD
4.2 Multimodal Open API (MOAPI) for Rich Hybrid Queries
To achieve our goal of conducting rich hybrid queries on MQRLD, a versatile and easy-to-use query interface is essential. SQL, as a wildly used query language, performs well in processing structured data and relational operations. However, it faces challenges when comes to concisely expressing complex query logic over multimodal data that cannot be defined by tables as entities and relationships. In contrast, open API provides versatile interface definitions, enabling it to handle querying a wide variety of data types. The Jina API (JINA, 2024) is a well-behaved open API that is specifically designed for interpreting and interacting with multimodal data, providing a uniform interface specification for querying multiple data types. Therefore, our MOAPI utilizes Jina API as the query interface and defines four basic query types, including:
  • Numeric Equal (N.E) Query: Returns results that are equal to the given value of an attribute.
  • Numeric Range (N.R) Query: Returns results where the value of an attribute falls within the given range.
  • Vector KNN (V.K) Query: Given a set $P$, a query object $q\in P$, and a positive integer $k$, returns $k$ objects in $P$, denoted as V.K($q$,$k$), such that $\forall p\in\text{V.K}(q,k),p^{\prime}\in P\backslash\text{V.K}(q,k),dist(q,p)\leq dist(q,p^{\prime})$.
  • Vector Range (V.R) Query: Given a set $P$, a query object $q\in P$, and a query radius $r\geq 0$, returns all objects in $P$ within the distance $r$ of $q$, i.e., V.R($q$,$r$)=$\{p\in P|dist(p,q)\leq r\}$.
[…]
4.3 Statistic Table for Query-aware Mechanism
We establish the query-aware mechanism by building a Query Behavior Statistic (QBS) table and recording a series of insightful statistic variables during the query execution. This QBS table and its statistic data play a vital role in constructing MQRLD through (1) reflecting the performance of query tasks to facilitate feature measurement process (Section 5.1.2), (2) identifying the feature of multimodal data representation in different query context to improve feature data layout (Section 5.2.2), and (3) dynamically adjusting and improving index structure to provide faster and more accurate retrieval (Section 6.2).
Table 3: Query Behavior Statistic Table.
Query Statement Query Multimodal Object Set Query Attributes Query Type Recall@K CBR Query Time Query Accuracy
Task1 DF1 Em1, Em2 NE,VK R1 X1 T1 A1
Task2 DF2 Em3 VR R2 X2 T2 A2
As shown in Table 3, the statistic variables recorded in the QBS table are as follows:
  • Query Statement: The input query statement.
  • Query Multimodal Object Set: The DataFrame involved multimodal objects in the query statement.
  • Query Attributes: The columns involved in the query statement. In hybrid queries, achieving accurate results often requires utilizing two data attributes, while in rich hybrid queries, we may introduce more data attributes into the query process.
  • Query Type: The four basic query types performed on each attribute.
  • Recall@K: The proportion of correct matches among the top K retrieved results.
  • CBR (Cross Bucket Rate): The evaluation metric for the index effectiveness, akin to the "Inter-bucket Traversal Rate" or "Cross-partition Traversal Rate". We utilize the DataFrames partitions to correspond to the "buckets" or "partitions".
  • Query Time: The average query execution time for queries of the specific task.
  • Query Accuracy: The ratio between the query results and the ground truth.
5 Multimodal Data Feature Representation
The goal of multimodal data feature representation is to obtain optimal features and data layouts for multimodal data. This process is based on insights from both the data itself and the queries performed on it, aiming to improve the quality of the high-dimensional learned index and ultimately improve query performance. As shown in Fig 5, this process can be divided into two parts: (1)Feature Embedding and Measurement: We obtain representative features of multimodal data using our unique measurement process by concerning intrinsic and extrinsic metrics to select the embedding model with the highest score so that the features can express multimodal data more effectively. (2)Feature Enhancement: We optimize features through hyperspace transformation and movement by analyzing data patterns and learning from query behaviors to adjust feature vectors and overall data layout, making index and rich hybrid queries more efficient.
Figure 5: Overview of Multimodal Data Representation
5.1 Feature Embedding and Measurement
Figure 6: Workflow of Feature Embedding and Measurement
Feature embedding is the process of converting multimodal raw data into a set of features, while the measurement procedure serves as an assessment to determine the effectiveness of the embedding model, aiding in the selection of the more representative features for subsequent indexing, and querying tasks. We first provide an overall workflow of embedding and measurement, then delve into details. The overall workflow is shown in Fig 6. In the feature embedding process, MQRLD offers an embedding model pool for embedding the given multimodal raw data. The embedded features are then subjected to the measurement process, where we score each model using a data-aware intrinsic metric and a query-aware extrinsic metric to select the embedding model with the highest score.
5.1.1 Embedding
To initialize, we introduce the construction of the embedding model pool. In the field of feature embedding domain, numerous existing works have proposed diverse techniques for efficiently representing raw data. Table 4 lists some of the notable works that have proven successful in multimodal data feature embedding, including image data, text data, audio data, etc., which we have collected and consolidated into the embedding model pool to support the embedding of wide-ranging attributes of multimodal data. Nevertheless, there is still a critical need to select a suitable model for different scenarios, as the performance of these embedding models can vary significantly depending on factors such as data characteristics and query requirements. For instance, compared to CLIP4Clip, X-CLIP uses temporal embedding models to capture dynamic actions and interactions over time, making it more suitable for the task of embedding features from a given soccer match video. Conversely, CLIP4Clip can effectively capture static, frame-level information, which is more appropriate for life recordings that consist of fragmented information with weak temporal correlations. Consequently, measurement metrics are needed to measure the embedding features, demonstrating the effectiveness of each embedding model.
Table 4: Pre-trained Cross-Modal Embedding Models in Embedding Model Pool.
Embedding Model Image Embedding Text Embedding Video Embedding Audio Embedding
CLIP (Radford et al., 2021)
CLIP4Clip (Luo et al., 2021)
X-CLIP (Ma et al., 2022)
AudioCLIP (Guzhov et al., 2022)
OPT (Liu et al., 2021)
5.1.2 Measurement
Evaluating feature embedding models for multimodal data is a challenging task because it requires consideration of generalization ability, fidelity, and support for high-quality downstream query tasks. To address this, we propose a combined measurement metric: widely-used extrinsic metric, complemented by our innovative intrinsic metric, to provide a coarse-to-fine-grained evaluation paradigm. As shown in Fig 6, we define a scoring system to evaluate the effectiveness of each modal, which is formulated as:
$Score=w_{1}S_{1}+w_{2}S_{2}+w_{3}S_{3}$ (1)
where the weights ${w_{1},w_{2},w_{3}}$ are used to control the influence of different metrics on the final measurement. $S_{1}$ is an extrinsic measurement metric, focusing on the performance in downstream tasks, while $S_{2}$ and $S_{3}$ are intrinsic measurement metrics, emphasizing generalization and fidelity ability, respectively.
To be specific, extrinsic metric $S_{1}$ is obtained after executing the specific downstream task by measuring three metrics: Recall@K, Query Accuracy, and Query Time, which are recorded in the QBS table during the query process. In contrast, intrinsic metrics $S_{2}$ and $S_{3}$ are obtained by analyzing the dataset itself, by calculating Silhouette Coefficient (SC) (Rousseeuw, 1987) and Fréchet distance (FID) (Heusel et al., 2017), respectively.
The remainder of this section will introduce these two metrics in detail, and finally present experimental validations demonstrating that our measurement pipeline improves the selection of feature embedding models.
Silhouette Coefficient (SC)
SC is measured from the perspective of the whole dataset. Features embedded by an effective embedding model from raw data should exhibit well-clustered characteristics, that is, the intra-class distance (distance within each class) should be small and the inter-class distance (distance between different classes) should be large, facilitating feature-related retrieval. To assess the quality of clusters formed by embedding models, we innovatively introduce the SC metric to provide a coarse-grained score that indicates both the consistency within clusters and the separation between clusters. Specifically, we embed the features using $n$ candidate embedding models:
$X_{n}=\{\text{Model}_{1}(I),\text{Model}_{2}(I),.....\text{Model}_{n}(I)\},$ (2)
where $I$ is the input data, and $X_{n}$ is the $n$ sets of features embedded from $n$ candidate models. Then we cluster each set of features into $k$ classes:
$\text{Cluster}(X_{n},k)=\{\{x_{1}^{i}\}_{i=0}^{k},\{x_{2}^{i}\}_{i=0}^{k},...,\{x_{n}^{i}\}_{i=0}^{k}\},$ (3)
where Cluster is a common clustering method, which can be K-means (Macqueen, 1967) or GMM (Reynolds et al., 2009), and $\{x_{j}^{i}\}_{i=0}^{k}$ is the clustered features encoded by $Model_{j}$. Next, we obtain the score of $S_{2}$ for each model by calculating the SC value of the clustered features:
Score of $\text{Score of }S_{2}\text{ for }Model_{j}=SC(\{x_{j}^{i}\}_{i=0}^{k}),$ (4)
where $SC()$ is the function to calculate the SC value of a given cluster results.
Fréchet distance (FID)
Unlike the global consideration of SC, FID focuses on feature quality for each individual multimodal object. Recent advancements in text-to-image diffusion models have yielded impressive results in generating realistic and diverse images, for example, the remarkable progress in Stable Diffusion makes it possible to generate high-fidelity images with a CLIP text embedding model. Thanks to CLIP, building a bridge between text and image, we could reconstruct the embedded image with a powerful diffusion model, which gives us a chance to have a comprehensive understanding of the visual essence. FID reflects the difference between the original image and the generated image. We calculate the FID of each image’s fine-grained feature distribution and use the FID to judge the fidelity of feature embedding for each multimodal object. Specifically, for embedding model $Model_{j}$, we input the embedded features into the same pre-trained high-performance generative model like Stable Diffusion (Rombach et al., 2022) to get the reconstructed images $\hat{I}_{j}^{{}^{\prime}}$, $\hat{I}_{j}^{{}^{\prime}}=\text{Diffusion}(X_{n})$. Then we input reconstructed images in turn with the original images $\hat{I}_{j}$ into the inception network to get FID value $\text{FID}_{j}=\text{Inception}(\hat{I}_{j},\hat{I}_{j}^{{}^{\prime}})$, and the score of $S_{3}$:
Score of $\text{Score of }S_{3}\text{ for }Model_{j}=1-\text{Normalization}(\text{FID}_{j})$ (5)
A small FID value between the generated image and the original image indicates that the embedding model has excellent performance and can effectively learn the key information from the original data.
Integrating both coarse-grained and fine-grained measurements with a weighted approach, we derive the final inner score (combination of $S_{2}$ and $S_{3}$). This inner score is then fused with the external score, allowing for a comprehensive selection of the embedding model for different scenarios.
Experimental Validations
To demonstrate the effectiveness of our measurement method, we select a set of image feature embedding models for evaluation. The experiments utilize a real dataset AI Challenger, with evaluation methods including the SC (Silhouette Coefficient), IN (Intrinsic Measurement), and IN + EX (a combination of Intrinsic Measurement and Extrinsic Measurement). The scoring results are calculated based on different evaluation methods:
${Score}=\left\{\begin{array}[]{ll}S_{2}&\text{ if method }=\mathrm{SC}\\ w_{2}S_{2}+w_{3}S_{3}&\text{ if method }=\mathrm{IN}\\ w_{1}S_{1}+w_{2}S_{2}+w_{3}S_{3}&\text{ if method }=\mathrm{IN}+\mathrm{EX}\end{array}\right.$ (6)
where the weights for the IN are $w_{2}=0.3$ and $w_{3}=0.7$. For the IN + EX, the weights are $w_{1}=0.2$, $w_{2}=0.3$ and $w_{3}=0.5$.
In the experiment, we select the image feature embedding models RN50, ViT-B/16, and RN50x64 as experimental benchmarks. Fig 7 illustrates the scoring result of different feature embedding models across various evaluation methods compared with the result of downstream query tasks. The x-axis represents different sizes of image samples and downstream query tasks in the experiment, and the y-axis represents the scoring results after normalization. The scoring results for the downstream query task show that RN50x64 achieves the highest score, followed by ViT-B/16 and RN50, reflecting the actual effectiveness of their respective features in the real query task after embedding. This scoring method effectively distinguishes the performance of different embedding models for downstream query tasks and aligns with the actual scoring results. Comparing the scoring results from different evaluation methods (SC, IN, IN + EX) with those of the downstream query tasks, we find that the SC scoring results show more generalization, as they always fluctuate in a small interval. On the other hand, scoring results IN and IN + EX show more variation and align closely with the scoring results of downstream query tasks. This differentiation arises because SC emphasizes the generalization ability of the embedding model, while IN integrates FID which measures the fidelity of image features, improving evaluation accuracy compared to SC. IN + EX evaluation incorporates feedback from downstream query tasks, providing a more accurate assessment of each embedding model and coming closer to their actual effectiveness.
Figure 7: Comparative Evaluation Results of Different Embedding Models
5.2 Feature Enhancement
A well-designed representation of multimodal data is crucial for reducing index scan numbers and speeding up queries. This leads to our feature enhancement process that effectively adjusts the feature vectors and data layout by gathering similar multimodal data together, which can improve the quality and speed of the high-dimensional index and ultimately enhance query performance. To be specific, as shown in Fig 8, we treat a high-dimensional feature vector as a data point in hyperspace, and then relocate it to an optimal new location through a unique "projection", considering both the dataset characteristics and query behaviors (Section 5.2.1). The specific method of this projection includes Hyperspace Transformation (Section 5.2.2) and Hyperspace Movement (Section 5.2.3). In the former, the matrix representing the hyperspace transformation is determined, while in the latter, we make further fine-tuning movement. This projection allows the most discriminative dimensions of the data to be emphasized, and similar feature vectors are clustered closely together. Then, we provide evidence that our method successfully improves the performance of clustering within the dataset (Section 5.2.4), which forms a solid foundation for index construction.
Figure 8: An Overview of Feature Enhancement
5.2.1 Awareness of Projection
There are two aspects we take into account in the projection process, which are the dataset characteristics and query behaviors. From the perspective of dataset, we have the following considerations: (1)Identify and retain information-rich dimensions of certain attributes; (2)Discern underlying distribution of entire dataset; (3)Utilize relationships between multimodal objects.
The first two considerations are applied during the hyperspace transformation phase in Section 5.2.2, while the third consideration is applied during the hyperspace movement phase in Section 5.2.3. From the perspective of query behaviors, different queries may focus on different features when searching for multimodal data. Therefore, we add an optimization step to the process of hyperspace transformation to facilitate locating relevant data.
5.2.2 Hyperspace Transformation
The hyperspace transformation converts all data points (corresponding to all feature vectors) into new, optimal positions by considering the characteristics of the data and the specific query context. This is achieved through a transformation matrix $T$, which can be derived and optimized in four steps, as detailed below:
Step1:  Represent data with matrix $D$
All the multimodal objects are managed in DataFrame, from where we select columns of DataFrame to be indexed and represent them with a matrix $D$, as shown in Fig 9, where $D\in\mathbb{R}^{m\times n}$, with each row in matrix $D$ corresponding to a record of a multimodal data object in the DataFrame. $m$ is the total number of records in the DataFrame, and $n$ is the dimension number of all the feature vectors in selected columns.
Figure 9: Matrix D
To enhance the features of multimodal data without further compressing the information and maintain traceability to the original dataset, we define an $N\times N$ transformation matrix $T$ that establishes a one-to-one mapping. Then, we multiply the original matrix $D$ by the $T$ to get enhanced feature vectors denoted as $D_{T}$, where $D_{T}\in\mathbb{R}^{m\times n}$, and $D_{T}=DT$.
Step2:  Compute the covariance matrix $C$
By analyzing matrix $D$, we can compute the covariance matrix $C$ ($C\in\mathbb{R}^{n\times n}$) of $D$. The covariance matrix $C$ not only reflects the joint distribution characteristics of the data but also reveals the importance of each dimensions in feature vectors. The $n$ rows of $C$ highlight the distribution pattern of all data records. If the rows of $C$ exhibit small fluctuations, their differences from the mean will be minimal, resulting in a small covariance value. The $n$ columns imply the importance of each dimension and also preserve all dimensional information of the selected columns in the DataFrame, ensuring the specific columns can be accurately mapped using the matrix.
Step3:  Define the transformation matrix $T$
Through the decomposition of $C$, we can construct the transformation matrix $T$ applied to the feature vectors. $T$ comprises a rotation matrix $R$ and a scaling matrix $S$, i.e., $T=RS$. The rotation matrix $R$ indicates the direction of data transformation in hyperspace, while $S$ is a diagonal matrix reflecting the importance of each vector dimension through the scaling of its diagonal elements, thus stretching each dimension of feature vectors. $R$ and $S$ are calculated though eigen decomposition on the covariance matrix $C$ which is represented as:
$C=V\Lambda V^{T}=\begin{bmatrix}v_{1},v_{2},\cdots,v_{n}\end{bmatrix}\begin{bmatrix}\lambda_{1}&0&\cdots&0\\ 0&\lambda_{2}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&\lambda_{n}\end{bmatrix}\begin{bmatrix}v_{1}^{T}\\ v_{2}^{T}\\ \vdots\\ v_{n}^{T}\end{bmatrix}$ (7)
where $V=(v_{1},v_{2},\cdots,v_{n})$ is a matrix that is composed of all the eigenvectors of $C$, and $\Lambda=diag(\lambda_{1},\lambda_{2},\cdots,\lambda_{n})$ is a diagonal matrix, with its diagonal elements being the eigenvalues corresponding to each dimension.
According to the linear transformation relationship of matrices, we know that the eigenvectors in $V$ serve as the basis for the linear transformation of hyperspace coordinates, and the square root of each eigenvalue in $\Lambda$ represents the scaling factor for each dimension in hyperspace. Therefore, we can obtain the rotation matrix $R=V$ and scaling matrix $S=\sqrt{\Lambda}$ from the eigen decomposition results of $C$, and the transformation matrix $T$ can be defined as:
$T=RS=\begin{bmatrix}v_{1},v_{2},\cdots,v_{n}\end{bmatrix}\begin{bmatrix}\sqrt{\lambda_{1}}&0&\cdots&0\\ 0&\sqrt{\lambda_{2}}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&\sqrt{\lambda_{n}}\end{bmatrix}$ (8)
The invertibility of the hyperspace transformation is an innegligible issue, because after performing indexing, the indexed data needs to be re-transformed into the original feature vector to support query operations on the original data. Therefore, to ensure the reversibility of matrix $T$ and maintain the linearity of data transformation properties, we impose the following constraints on $R$ and $S$:
$(1)R,S\in\mathbb{R}^{n\times n}.$ (9)
$(2)\forall v_{i},v_{j}\in R,||v_{i}||_{2}=1,v_{i}^{T}v_{j}=0,\forall i,j\in\{1,2,\ldots,n\},i\neq j.$
$(3)S=diag(a_{11},a_{22},\cdots,a_{nn}),a_{ii}>0,\forall i\in\{1,2,\ldots,n\}.$
Constraint (1) indicates that both the rotation matrix $R$ and the scaling matrix $S$ are $n\times n$ square matrices, ensuring that matrix $T$ remains a $n\times n$ square matrix throughout multiply of $R$ and $S$. As a result, after the transformation, the dataset retains $n$ dimensions in the hyperspace with no loss of dimensional information, ensuring that the original dataset is easily regained through one-to-one mapping. Constraint (2) guarantees that $R$ is an orthonormal matrix, with its column vectors forming the standard orthonormal basis of the hyperspace. This ensures that the columns of the matrix transformed by $R$ are independent of other columns, allowing us to focus on specific dimensions corresponding to the feature vectors during query-aware optimization without affecting other dimensions. Constraint (3) specifies that matrix $S$ must be a positive definite matrix, because we only consider scaling the feature vectors in the positive direction. By then, $T$ is calculated and we can get enhanced feature vectors $D_{T}$, where $D_{T}=DT$. These three constraints ensure that matrix $T$ is invertible, which allows us to obtain the original feature vectors directly through the inverse matrix operation by multiplying $D_{T}$ by $T^{-1}$.
Step4:  Optimize $T$ by query awareness
We consider the specific requirements of query tasks and adjust the feature vectors to better align with the query context, thereby improving query efficiency and accuracy. To ensure that feature representations accommodate different query contexts, we adjust the hyperspace transformation process by optimizing $T$ through learning from query behaviors. To simplify this optimization process, we decompose the objective of optimizing $T$ into optimizing $R$ and $S$.
To be specific, our goal is to find the optimal $R^{*}$, $S^{*}$ that maximizes the query performance of query workload $Q$, i.e., we formulate this as an optimization problem as
$R^{*},S^{*}=argmax_{Q}QueryScore(R,S)$ (10)
where the $QueryScore$ function returns a scalar score value that evaluates the performance of query execution, with the assumption that a higher score value indicates better performance.
When modeling the $QueryScore$ function, we consider three factors, which are query accuracy, query time, and CBR. This is based on the intuitive idea that an effective index leads to more accurate queries, fewer cross-buckets (resulting in lower I/O), and shorter query times. The detailed description of these three metrics is in Section 4.3.
Based on this consideration, we model $QueryScore$ function:
$QueryScore=\alpha_{1}Score_{acc}-\alpha_{2}Score_{CBR}-\alpha_{3}Score_{time}$ (11)
where $Score_{acc},Score_{CBR},Score_{time}$ are the query scores corresponding to the above metrics of query accuracy, CBR, and query time. The weight value $\alpha=\{\alpha_{1},\alpha_{2},\alpha_{3}\}$ is set to standardize each score value and eliminate the impact of different magnitudes among variables.
Given a query workload $Q$, it would be prohibitively costly to directly compute the individual score for each metric. Therefore, we train a random forest model to approximate each score value:
$(Score_{acc},Score_{CBR},Score_{time})=(f_{acc}(R,S),f_{CBR}(R,S),f_{time}(R,S))$ (12)
where $f_{acc}(R,S),f_{CBR}(R,S),f_{time}(R,S)$ are three random forest functions that are used to estimate three score values, which are trained with $R$, $S$, and statistical data from QBS table.
According to equation 10, our goal is to find the optimized matrix $R^{*}$ and $S^{*}$ that maximize the $QueryScore$. The optimization process is presented in Algorithm 1 based on SMAC3 (Lindauer et al., 2022). Matrix $R$ and $S$ are initialized from the original dataset $D$ as described in Step 3, and surrogate model $M$ is obtained using equation 12. Then we iteratively update $S$ using the EI acquisition function and reconstruct $R$ based on updated $S$ to calculate new score value. After $k$ iterations, we return $R^{*}$ and $S^{*}$ that have the highest score value. In this way, we can get optimized feature vectors $D_{T^{*}}=DT^{*}$, where $T^{*}=R^{*}S^{*}$.
0:  dataset matrix $D$, workload $Q$, max iteration k, surrogate model $M$
0:  Optimized matrix $R^{*},S^{*}$
1:  Initialize $R$ and $S$ based on $D$, $Score_{best}=0$, $R_{best}=R$, $S_{best}=S$
2:  for i in [0,k] do
3:     $S_{next}=SelectNext(M,S_{best})$ /* use EI acquisition function to select new sample points on the surrogate model $M$*/
4:     $R_{next}=ReConstruct(S_{next})$
5:     $Score_{next}=BatchEval(S_{next},D,Q)$
6:     Update $M$ based on $(S_{next},Score_{next})$
[…]
8:        $Score_{best}=\text{max}(Score_{next})$
9:        $S_{best}=argmax(Score_{next})$
10:        $R_{best}=R_{next}$
11:     end if
12:  end for
13:  return  $R_{best},S_{best}$
Algorithm 1 Matrix Optimization
5.2.3 Hyperspace Movement
Hyperspace movement further optimizes the feature vectors on the result of previous hyperspace transformation by moving similar data points in hyperspace together to get $\hat{D}$, where similar data can be placed in adjacent places or the same cluster. Thus it could benefit index building and multimodal data retrieval.
Figure 10: Hyperspace Movement
To achieve this objective, we propose a new data movement method called Local Parallelized Gravitational Field (LPGF) which improves on an existing data movement method HIBOG (Li et al., 2021). HIBOG posits that in hyperspace, a data point is attracted by its similar data points, causing it to move towards the resultant force direction of these attractions. Specifically, it defines a gravitation force between each data point and its nearest $K$ points and applies this force to this data point to make displacement. Though HIBOG shows effectiveness in gathering similar data, it suffers the following drawbacks: (1) Computationally expensive: Finding the nearest $K$ data points for force calculation requires sorting all data points, which requires huge computation, especially for large datasets. Additionally, the parameter $K$ needs to be set and adjusted, further increasing the computation time. (2) Anomalies of movement: The calculation of forces between data points may not be suitable for all scenarios. Sometimes, this force can cause anomalies in the movement of tightly clustered data. (3) Weak parallel capability: The spatial range of the nearest $K$ neighbors is not predetermined, leading to the inability to perform grid partitioning and parallel processing of data. To overcome these shortcomings, we introduce the pipeline of LPGF, which consists of three steps to improve the above shortcomings, as shown in Fig 10.
Step1:  Boundary of force area
[…]
Step2:  Distinct regions of force
To avoid anomalies of movement, we make the strength and direction of applied force on point $\overrightarrow{P_{i}}$ depend on the distance between $\overrightarrow{P_{i}}$ and all the other points within radius $R$. To be specific, for each $\overrightarrow{P_{i}}$, the resultant force $\overrightarrow{F_{i}}$ is calculated by summing the applied force from all relevant points $\overrightarrow{P_{i}}$ in $R$. $\overrightarrow{F_{ij}}$ applied by $\overrightarrow{P_{ij}}$ depends on the distance between $\overrightarrow{P_{i}}$ and $\overrightarrow{P_{ij}}$, which is calculated as Fig 11, where $C$ is a constant slightly larger than 1. For example, taking a value of $1+10^{-1}$. During the calculation of gravitational forces, data points are represented in vector form. Within the radius $R$ of data point $\overrightarrow{P_{i}}$, the resultant force $\overrightarrow{F_{i}}$ exerted on $\overrightarrow{P_{i}}$ is determined by $\overrightarrow{F_{i}}=\sum_{j=1}^{N}\overrightarrow{F_{ij}}$, where N is the number of data points contained within the radius $R$ of $\overrightarrow{P_{i}}$.
Figure 11: Force Calculation Formula
Step3:  Parallelization of force computation
To improve parallel capability, we split the dataset into multiple regions and utilize the partitioning mechanism of Apache Hudi to perform the computation in parallel within each region. This is achievable because we define a padding area with radius $R$ to limit the range for processing each point, ensuring that the calculation area for each data point is confined to a fixed region. After applying LPGF, we can obtain a displacement matrix $M$, $M\in\mathbb{R}^{m\times n}$, which represents the displacement of each data point.
5.2.4 Feature Representation Evaluation
To prove that our method could successfully improve the performance of clustering within the dataset, which forms a solid foundation for index construction, we design experiments to demonstrate the improvement of clustering (Evaluation One) and high-dimensional learned index query performance (Evaluation Two) by our proposed feature representation method.
Evaluation One:  Clustering Enhancement Evaluation
To validate the capability of our proposed methods to gather similar multimodal data together, we compare five feature representation methods for enhancing clustering performance, which are T, HIBOG, LPGF, T + HIBOG, and T + LPGF. T represents Hyperspace Transformation. HIBOG and LPGF denote Hyperspace Movement. T + HIBOG as well as T + LPGF represent the combination of Hyperspace Transformation and Hyperspace Movement. The experiments are conducted using the real dataset Flame (Fu and Medico, 2007) and employ three well-known clustering methods, which are K-means, Agglomerative (Sokal and Michener, 1958), and Density Peaks Clustering (DPC) (Rodriguez and Laio, 2014), to represent diverse clustering paradigms. The selection of these methods is not meant to imply exclusivity but rather to demonstrate the generalization ability across various clustering algorithms. To comprehensively evaluate the impact of different representation methods on clustering performance, we use the Silhouette Coefficient, Calinski-Harabasz Index, and Normalized Mutual Information as experimental metrics. These metrics assess clustering effectiveness from different perspectives.
Table 5: Clustering Enhancement Results of Feature Representation
Clustering Method Optimization Method Silhouette Coefficient Calinski-Harabasz Index Normalized Mutual Information
K-means Unoptimized 0.412 202.695 0.476
T 0.425 214.568 0.506
HIBOG 0.448 222.804 0.523
LPGF 0.455 228.109 0.541
T + HIBOG 0.472 245.784 0.576
T + LPGF 0.503 260.556 0.603
Agglomerative Unoptimized 0.329 122.782 0.330
T 0.342 129.652 0.377
HIBOG 0.363 144.331 0.425
LPGF 0.364 144.868 0.425
T + HIBOG 0.395 151.546 0.489
T + LPGF 0.423 159.651 0.541
DPC Unoptimized 0.338 133.615 0.413
T 0.345 135.781 0.614
HIBOG 0.351 138.855 0.890
LPGF 0.410 169.101 0.936
T + HIBOG 0.455 176.252 0.958
T + LPGF 0.491 184.451 0.992
  • a
    Bold font indicates the best result, and italic font indicates the second-best result.
The experimental results of different feature representation methods are presented in Table 5. Comparing the results of LPGF and HIBOG, we observe that LPGF significantly enhances clustering performance across various algorithms, suggesting that optimizations in gravitational field design in LPGF improve intra-class compactness and inter-class separation more effectively than HIBOG. Additionally, T + HIBOG and T + LPGF demonstrate that the hyperspace transformation matrix T not only enhances the discriminative features of multimodal data but also substantially improves clustering effectiveness.
Evaluation Two:  High-dimensional Learned Index Query Performance Enhancement Evaluation
We design two experiments to demonstrate the effectiveness of our proposed feature representation method in enhancing the performance of high-dimensional learned index queries. The first experiment demonstrates that our method can achieve excellent "last-mile" models within the leaf nodes, while the second evaluates their influence on query performance of currently typical high-dimensional learned indexes. The experiments utilize a generated dataset comprising 2M records, with data distributed according to a Gaussian mixture distribution.
In the first experiment, we compare the smoothness of Cumulative Distribution Function (CDF) curves of datasets processed by different feature representation methods. Smooth CDF curves help address the "last mile problem" in learned index, making it easier for simple machine learning and improving the accuracy of index predictions. We compare the CDF curves of four datasets, which are Original Dataset, HIBOG Optimized Dataset, LPGF Optimized Dataset and T + LPGF Optimized Dataset. As shown in Fig 12, by comparing the smoothness of the CDF curves for each dataset, we find that T + LPGF achieves the highest improvement in the smoothness among all optimized methods. Compared to HIBOG, LPGF significantly smoothens the global CDF curve, because the force calculation method in LPGF reduces the probability of anomalous data movement, resulting in tighter data clustering. T + LPGF not only retains the advantages of LPGF but also highlights critical clustering features through matrix T, further enhancing the clustering capabilities. Overall, the T + LPGF method results in tighter multimodal data clustering, improving the smoothness of the CDF curve, and effectively reducing the complexity of the machine learning involved in index construction.
Figure 12: The CDF Curves of Different Feature Representation Methods
Figure 13: The Average Query Times of Different Indexes on Original Dataset and Optimized Dataset
In the second experiment, we demonstrate the effectiveness of our methods in improving the efficiency of typical high-dimensional learned indexes by comparing the average query times. We select five typical high-dimensional learned indexes, which are ML, ZM, LISA, Flood, and LIMS. These learned indexes all share the characteristic that their query performance is significantly influenced by the smoothness of the CDF curve which directly impacts "last-mile" searching. As shown in Fig 13, we compare the average query times of different indexes on the original, the LPGF optimized, and the T + LPGF optimized datasets. The experimental results indicate that both LPGF and T + LPGF methods significantly improve the query performance, with T + LPGF performing the best. ML and LIMS show the most significant performance improvements across all datasets, because the creation of these indexes relies heavily on clustering results, confirming the effectiveness of the T + LPGF method in enhancing data clustering. LISA and Flood, being variants of grid indexes, benefit from the ability of T + LPGF to make data more tightly clustered, thereby improving the smoothness of the CDF curve within the grid. Since the calculation of z-addresses in ZM is closely related to the positions of data in hyperspace, a tighter data distribution can improve its continuity, making the data easier for machine learning. The T + LPGF method improves the overall distribution and continuity of data, providing smoother CDF curves that are easier for various indexes to learn, effectively enhancing the query performance of high-dimensional learned indexes.
6 High-dimensional Learned Index
To fully leverage the advantages of the "Multimodal Data Feature Representation" introduced in Section 5 and further enhance multimodal data retrieval efficiency, we propose a high-dimensional learned index. This index hierarchically organizes clusters and their sub-clusters to form a cluster tree, where the index of the leaf nodes is established through a "last-mile" training approach. Furthermore, the index’s inner structure can be optimized through our query-aware mechanism. An overview of our index is illustrated in Fig 14, detailing both the construction and optimization processes. During the construction phase, divisive hierarchical clustering is employed to form clusters at various levels, corresponding to different levels within the tree structure, with the leaf nodes representing the most granular, indivisible sub-clusters, thereby substantially enhancing the precision of the "last-mile" training for the index. Concurrently, the optimization phase employs a query-aware mechanism to refine the internal structure of the index, improving its adaptability to diverse query demands. The detailed processes of index construction and optimization are elaborated in Section 6.1 and Section 6.2, respectively.
Figure 14: Overview of Index Construction
6.1 Index Construction
The process of constructing the index comprises two primary stages, which are divisive hierarchical clustering with "last-mile" training and cluster tree construction.
6.1.1 Divisive Hierarchical Clustering with "Last-mile" Training
This stage splits data points through recursive partitioning, creating nested clusters that progressively divide the dataset into smaller and more refined subsets, delineating the hierarchical level of each subset. This nested clustering approach enables a hierarchical narrowing of the search scope, followed by "last-mile" training in the most granular subsets for faster lookup. The entire Divisive Hierarchical Clustering algorithm can be described as an iterative step of "Division" and "Training based Evaluation", as outlined in Algorithm 2.
Step 1:  Division
In the division step, the intermediate result set from the previous iteration is further divided into smaller subsets, while the training-based evaluation step assesses whether the subsets obtained from the previous iteration have reached an optimal distribution. If optimal distribution is reached, the division stops, and the training results for each optimal subset are saved for the "last-mile" search. Otherwise, the division continues. The iteration stops when all subsets from the previous iteration have achieved an optimal distribution, resulting in the final clustering hierarchy and completing the "last-mile" training for all minimal subsets.
We begin by treating the entire dataset $\hat{D}$, processed through the previous multimodal data feature representation process, as a single cluster. Using the DPC algorithm, we then partition this dataset into initial sub-clusters $\{\hat{D_{i}}\}$.
The selection of the DPC algorithm is motivated by its ability to automatically ascertain the optimal number of sub-clusters, thus eliminating the necessity for manual intervention. This capability is particularly beneficial in contexts where the exact number of clusters is not predefined. More importantly, DPC utilizes the results of LPGF where data points are more densely grouped, enhancing the clustering outcomes and ensuring more uniform data distribution within subsets.
Step 2:  Training based Evaluation
Following the step of obtaining initial sub-clusters $\{\hat{D_{i}}\}$, we assess whether the data points in each sub-cluster are well-distributed to support accurate "last-mile" training and the corresponding efficient searching, thereby determining whether further subdivision of the cluster is necessary. We represent the centroid of $\hat{D_{i}}$ as $C_{i}$ using the Average Mass method (Protter and Morrey, 1970), and calculate the distance from each point $p$ in $\hat{D_{i}}$ to $C_{i}$. If the points in $\hat{D_{i}}$ are sufficiently uniform, distances of all points in it should be evenly spread, so that a simple linear regression model is sufficient to fit their distribution. Therefore, we predict the search position $v$ of a given lookup point $p$ using a model $v=F(p)*|\hat{D_{i}}|$, where $F(p)$ is the estimated CDF obtained through linear regression, which estimates the likelihood that a point’s distance to $C_{i}$ is less than or equal to the distance from the lookup point $p$. The accuracy of this model prediction validates the goodness that the CDF fits and further reflects the distribution of the points in $\{\hat{D_{i}}\}$. Therefore, we want the hit ratio of prediction for all points in the cluster to be greater than a given threshold $\delta$. Let $v(p)$ be the predicted searching position of point $p$ whose actual position is $v_{a}(p)$, our requirement is formulated as:
$\frac{1}{|\hat{D_{i}}|}\sum_{p\in\hat{D_{i}}}{IsEqual(v_{a}(p),v(p))}\geq\delta$ (13)
where the parameter $\delta$ is set to the optimal value 95.1% based on tuning with considerations of tree depth and query time (see Fig 15(a)). If this requirement is fulfilled, the points in $\hat{D_{i}}$ are uniform enough and do not require further partitioning. In this scenario, the model $F(p)$ will be stored for predicting the search position within $\hat{D_{i}}$. Otherwise, $\hat{D_{i}}$ must be further subdivided to achieve a better data distribution within its sub-clusters.
Figure 15: Calculation of $\delta$(a), Depth of Index Tree(b) and Leaf Node Characteristics for Different Dataset Sizes(c).
Through experimental verification, we set $\delta$ to 95.1%. The experiments also demonstrate that the depth of our tree does not significantly increase with the rapid growth of data volume. Detailed results are shown in Fig 15(a)-(c).
Fig 15(a) shows the cluster tree depth (we set the depth of the root node to 0) and average query time for various $\delta$ values. It can be observed that the best average query performance is achieved when $\delta$=95.1%. Although smaller $\delta$ values can further reduce the cluster tree depth, they lead to larger leaf nodes, requiring the scanning of more irrelevant data to obtain query results. Conversely, higher $\delta$ values increase the cluster tree depth, resulting in more node traversals during the query process.
Fig 15(b) presents experimental results on the depth of the cluster tree for Uniform, GuassMix, and Skewed datasets with different dataset sizes. The depth of the cluster tree increases gradually with the data volume, but the overall growth trend of the tree layers becomes more gradual. From the perspective of data distribution, the Skewed dataset results in the deepest cluster tree, reaching five layers. The maximum depths of the cluster trees for the GuassMix and Uniform datasets are the same, but the average depth of the cluster tree for GuassMix is significantly greater than that for Uniform. Though data skewness causes some tree branches to be more deeply divided, the depth does not grow excessively using our optimized feature representation strategies.
Fig 15(c) shows the number of leaf nodes, and the average data volume in the leaf nodes for Uniform, GuassMix, and Skewed datasets with different dataset sizes. As the data volume increases, both the number of leaf nodes and the average data volume within the leaf nodes increase. Observing the overall trend in the growth of leaf nodes, the data volume within each leaf node increases significantly, while the number of leaf nodes does not increase significantly. This is because the data-aware partitioning strategy controls node splitting based on the average error value, which is independent of the data volume.
0:  dataset $\hat{D}$ to be divided
0:  an array $A$ containing all subsets of $S$
1:  Initialize a queue of datasets $DatasetQueue=\{\}$
2:  $\{\hat{D_{i}}\}=DPC(\hat{D})$, $DatasetQueue.push(\{\hat{D_{i}}.resort()\})$
3:  while  $!DatasetQueue.isEmpty()$ do
4:     $S=DatasetQueue.pop()$
5:     $\{\hat{S}\},\{\hat{C}\}=DPC(LPGF(S))$ /*$\{\hat{S}\}$ is set of sub-clusters of S, and $\{\hat{C}\}$ is set of centroids of $\{\hat{S}\}$*/
6:     $C_{p}=getCentroid(\{\hat{C}\})$ /*computing parent cluster S’s centroid using sub-clusters’ centroid set $\{\hat{C}\}$*/
7:     $\{\hat{S}^{\prime}\},\{\hat{C}^{\prime}\}=resort(\{\hat{S}\},\{\hat{C}\},C_{p})$
8:     for $\hat{S_{i}}$ in $\{\hat{S}^{\prime}\}$ do
9:        $\{k\}=dist(C_{i},p\in\hat{S_{i}})$
10:        construct position prediction model $v_{i}=F_{i}(k)*|\hat{S_{i}}|$, where $F_{i}$ is the CDF model obtained from linear regression, formulated as $F_{i}(k)=a*k+b$
11:        $A.add(\hat{S_{i}})$
12:        if $\frac{1}{|\hat{S_{i}}|}\sum_{p\in\hat{S_{i}}}{IsEqual(v_{a}(p),v(p))}\leq\delta$ then
13:           $DatasetQueue.push(\hat{S_{i}})$
14:        end if
15:     end for
16:  end while
17:  return  $A$
Algorithm 2 Divisive Hierarchical Clustering
6.1.2 Cluster Tree construction
Following the divisive hierarchical clustering stage, the dataset is partitioned into nested sub-clusters, establishing a hierarchical structure of data points. We construct a cluster tree based on this partition result, encapsulating the order of subsets at each hierarchical level and the organization of each subset. At the top of this tree is the root node, representing the unique cluster that contains the entire dataset. As we traverse downward through the tree, each internal node represents a cluster that has been subdivided into smaller sub-clusters. An internal node with $n$ children signifies that its corresponding cluster has been partitioned into $n$ distinct sub-clusters. Furthermore, for each parent cluster $D_{p}$ with $n$ sub-clusters $\{D_{s}\}$, we calculate $D_{p}$’s centroid $C_{p}$ by Average Mass method using $\{C_{s}\}$, which is a set of centroids for $\{D_{s}\}$. The sub-clusters $\{D_{s}\}$ are then sorted based on the distances from their centroids $\{C_{s}\}$ to $C_{p}$ in increased order. This sorting is crucial for determining the search order among these sub-clusters, which is an important aspect when constructing the cluster tree. The leaf nodes represent the final, and undivided subsets. This hierarchical organization enables efficient data retrieval by progressively narrowing down the search space from the root to the pertinent leaf nodes.
To efficiently traverse the cluster tree, each node needs to store specific information that facilitates data searches within the node or navigation to a child. Based on this requirement, we record a tuple $t$ for each node. For non-leaf nodes, $t=\{C,R,L\}$, where $C$ is the centroid of the cluster corresponding to the node, and $R$ represents the cluster’s radius, defined as the Euclidean distance from $C$ to the farthest point in the cluster. Together, $C$ and $R$ define the search range for the data points within the node. $L$ is a pointer that directs to a list of its child nodes where these nodes are sorted and can be scanned based on the spatial order based on the sub-clusters’ distances to the parent node’s centroid. For leaf nodes, $t=\{C,R,M\}$. Again, $C$ and $R$ limit the search range of each leaf node. $M$ points to the search model trained in the previous step, which is used to predict the search position of an input lookup point, thereby facilitating rapid query responses.
0:  initial tree node list $L_{i}$, query workload $Q$
0:  optimized bucket order $\hat{L_{i}}$
1:  /*$B$ is an array of scanned times of each node after executing query workload $Q$, $t$ is query time*/
2:  $B,t=ExecuteQuery(Q,L_{i})$
3:  $\hat{L_{i}}$=sortInDescendingOrder($L_{i}$, B)
4:  startIter=0, endIter=0
5:  while endIter<|$\hat{L_{i}}$do
6:     size=0
7:     while endIter<|$\hat{L_{i}}$| and B[startIter]==B[endIter] do
8:        size+=1;endIter+=1;
9:     end while
10:     if size>1 then
11:        $L_{sub}=Permutations(\hat{L_{i}},\text{startIter},size-1)$
12:        for $l$ in $L_{sub}$ do
13:           $L_{l}=Reconstruct(\hat{L_{i}},\text{startIter},size-1)$
14:           $t_{l}=ExecuteQuery(Q,L_{l})$
15:           if $t_{l}$<$t$ then
16:              $t=t_{l}$,$\hat{L_{i}}=L_{l}$
17:           end if
18:        end for
19:     end if
20:     startIter=endIter
21:  end while
22:  return  $\hat{L_{i}}$
Algorithm 3 Tree Node Ordering Optimization
6.2 Index Optimization
Regarding query efficiency, different index structures are suited to different query workloads. In this section, we focus on optimizing and reordering sibling nodes (sub-clusters) that share the same parent node (parent cluster) based on the query-aware mechanism. This optimization aims to refine the index structure, thereby improving query performance and efficiency. Note that in our optimization strategy, we focus exclusively on the arrangement of all child nodes under each parent node while keeping the inheritance relationships unchanged. For instance, consider the scenario in Fig 14 where there are three non-leaf nodes ($N_{1}$,$N_{4}$,$N_{5}$) in the left branch of the cluster tree, each pointing to a list of their respective child nodes: $L_{1}=\{N_{4},N_{5}\}$, $L_{4}=\{N_{13},N_{14},N_{15}\}$, $L_{5}=\{N_{16},N_{17},N_{18}\}$. Our method is to optimize the order of nodes within each list (e.g. $L_{1}$ can be reordered as $L_{1}=\{N_{5},N_{4}\}$), without altering the parent-child relationships between nodes (e.g. $N_{5}$ cannot be moved to $L_{4}$).
This optimization is based on our observation that query patterns can vary with different query workloads, leading to certain nodes being accessed more frequently than others. By reordering the sibling nodes based on their access frequency, the index can be made more efficient. For example, as illustrated in Fig 14, if our lookup point ultimately resides in $N_{18}$, with the initial tree structure six nodes are required for scanning, while after optimization, the number of nodes scanned could be reduced to three, significantly saving query time.
Now we delve into the details of our optimization algorithm. Each non-leaf node $N_{i}$ stores a pointer to a sorted list which we denote as $L_{i}$. $L_{i}$ comprises all child nodes of the parent node $N_{i}$, initially ordered by their distance to the parent node’s centroid. We iteratively traverse all non-leaf nodes and, in $i^{th}$ iteration, our optimization goal is to adjust the ordering of $N_{i}$’s child node list $L_{i}$ to obtain a new list $\hat{L_{i}}$, so that the query time of a given query workload $Q$ is minimized:
$\hat{L_{i}}=argmin\sum_{q\in Q}{QueryTime(q,L_{i})}$ (14)
The optimization algorithm is shown in Algorithm 3. Its key idea is to move frequently accessed nodes to the beginning of the list to reduce the number of cross-leaf scans needed to find these "hot" nodes. We record the visit frequency of each node and sort them in descending order, placing the most frequently visited bucket at the head of the list. For nodes with the same visit frequency, we employ a brute-force approach to test each possible order and select the one that results in the minimum query time.
7 Experiments
We implement our MQRLD index in Scala. All experiments are completed in a distributed cluster environment built on the Ubuntu Linux system, using Apache Spark and Apache Hudi as the big data computing engine and data lake management framework, respectively. The cluster consists of three nodes, each with an Intel(R) Core(TM) i7-11700F CPU and 16GB memory.
7.1 Experimental Setup
7.1.1 DataSets
Table 6 shows the seven real-world datasets with different characteristics and three synthetic datasets we used in our experiments. The real-world datasets are also used in the previous work.
Table 6: Summary of Datasets.
Datasets Type Cardinality Dim. Size(GB)
OSM real 105M 6 5.04
Taxi real 184M 5 11.8
Stocks real 210M 4 13.2
Yelp real 6.99M 64 13.2
Color real 1.28M 32 4.2
Forest real 0.56M 12 1.5
AI Challenger real 0.30M N.A. 38
Uniform synthetic 1-100M 3-16 N.A.
GuassMix synthetic 1-100M 3-16 N.A.
Skewed synthetic 1-100M 3-16 N.A.
OSM [^1] https://download.geofabrik.de is a spatial dataset consisting of 105 million records randomly sampled from North America in the OpenStreetMap dataset. Each record includes six attributes, such as ID, timestamp, GPS coordinates, etc. Taxi [^2] https://www1.nyc.gov/site/tlc/about/tlc- trip- record- data.page is randomly sampled from records of yellow taxi trips in New York City in 2018 and 2019. We select five attributes from the dataset: the number of passengers, trip distance, pickup, drop-off locations, and total fare, to form a five-dimensional dataset. Stock [^3] https://www.kaggle.com/datasets/ehallmar/daily-historical-stock-prices-1970-2018 contains 210 million records, consisting of daily historical stock price data for over 6,000 stocks from 1970 to 2018. We focus on four features: trading volume, opening price, high price, and adjusted closing price. Yelp [^4] https://www.yelp.com/dataset consists of 6,990,280 user reviews of business places (e.g., cafes, restaurants, hotels, shops) from 11 metropolitan areas. For each review, we extract a set of 64-dimensional semantic vectors as the experimental dataset. Color [^5] https://image-net.org/download-images is a 32-dimensional image feature dataset extracted from the ImageNet dataset, which contains 1,281,167 records. Forest [^6] https://www.kaggle.com/c/forest-cover-type-prediction/data is collected by the US Geological Survey and the US Forest Service, which includes 565,892 12-dimensional records. AI Challenger [^7] https://challenger.ai is a multimodal dataset that contains more than 300,000 images with text descriptions. In order to simulate the rich hybrid queries scenario, we define two numeric attributes for each image: the length of the text description and the size of the image file. Synthetic Datasets include datasets with three different distributions, which are Uniform, GuassMix, and Skewed. The dimensions and data volumes of these datasets are adjustable. By default, we use a 3-dimensional dataset with a volume of 10 million records.
7.1.2 Query Generation
In the experiments, we assess the performance of range queries using the OSM, Taxi, and Stock datasets, and KNN queries using the Yelp, Color, and Forest datasets. Additionally, we evaluate the performance of rich hybrid queries based on the four basic query types proposed in Section 4.2, using the AI Challenger dataset.
The range queries and KNN queries, as common query operations, are each constituted by a record and its corresponding parameter, with the records being randomly selected from the dataset intended for querying. For range queries, the parameter is the query selectivity (i.e., the percentage of records that the query is expected to cover), which is set to a default of 10%, with a range between 0.1% and 20%. For KNN queries, the parameter is the number of query results $K$, which defaults to 1,000, ranging from 1 to 10,000.
For rich hybrid queries, we select three typical types of query among the rich hybrid queries shown in Fig 4, which are V.R $\bigoplus$ V.K, V.R $\bigoplus$ N.R and N.R $\bigoplus$ V.K. To verify the ability of MQRLD to cope with more complex queries, we also involve a query type of V.R $\times$ N, which is a combination of N instances of V.R (N $\in$ [2,5]). These queries are all generated from the AI Challenger dataset, whose image and text data can be embedded in multiple vector attributes for performing V.R or V.K queries. The parameters involved in rich hybrid queries include the query radius $R$ for V.R query, which defaults to 4% with a range from 0.1% to 10%; the parameter $K$ for V.K query, which defaults to 100 with a range from 1 to 1,000; and the selectivity for N.R query, which defaults to 10% with a range from 1% to 20%.
7.1.3 Competitors
We compare our index with other high-dimensional indexing methods, as shown in Table 7. All methods store data using DataFrame structures with data buckets as the smallest physical storage unit and apply the same optimizations where applicable.
Table 7: High-dimensional Index Competitors.
Index Type Range KNN
ZM (Wang et al., 2019) learned
ML (Davitkova et al., 2020) learned
LISA (Tian et al., 2022) learned
Qd-tree (Yang et al., 2020) learned
LIMS (Tian et al., 2022) learned
Flood (Nathan et al., 2020) learned
Tsunami (Ding et al., 2020) learned
R*-tree (Beckmann et al., 1990) tree
SPB-tree (Chen et al., 2015) tree
M-tree (Ciaccia et al., 1997) tree
7.1.4 Evaluation metrics
To evaluate the performance of the retrieval, we adopt the following two key metrics.
(1) Query time. To more accurately assess the performance of each index, we perform five repeated experiments for each group of generated queries and take the average query time as the experimental result.
(2) CBR. We use CBR to measure the dispersion of data access during the query process. A detailed definition refers to Section 4.3. A lower CBR indicates that the query is more concentrated within fewer data buckets, implying higher query efficiency and better indexing performance.
7.2 Evaluation of Range Query
In this section, we compare MQRLD with our competitors for multi-dimensional range queries. Fig 16 shows the query time for each optimized index on the OSM, Taxi, and Stocks dataset. MQRLD uses the multimodal data representation and high-dimensional learned index construction process in Section 5 and Section 6, while we tune the other methods as much as possible (e.g., ordered dimensions by selectivity and tuned the page sizes). This represents the best-case scenario for the other indexes, that the database administrator had the time and ability to tune the index parameters. To ensure the generalizability of our experimental results, we test the average query time of various index structures under different query selectivities. Our observations indicate that while the average query time for all indexes increases with higher query selectivity, MQRLD consistently outperforms other index structures across different datasets:
(1) Compared to tree-based indexes (R*-tree and Qd-tree), MQRLD demonstrates a significant improvement in query efficiency, with average performance gains of 8.1× and 3.9×, respectively. This is attributed to MQRLD’s ability to better capture the data distribution in high-dimensional space during its construction, retaining the advantages of tree-based indexes while ensuring similar data points are stored in the same or neighboring nodes.
(2) Against ZM, ML, and LISA, MQRLD shows average query efficiency improvements of 5.8×, 5.1× and 3.1×, respectively. A common feature of MQRLD and these indexes is the use of one-dimensional value mapping to expedite data lookup. For instance, LISA uses a mapping function based on data distribution to map similar data into the same segment. However, our data representation strategy goes a step further by clustering similar data based on their characteristics, resulting in a more efficient index structure.
(3) MQRLD achieves 2.4× and 3.3× speedup on query time compared to Flood and Tsunami, respectively. Although Flood’s cell division and Tsunami’s grid augmentation strategy can learn from query workloads to adjust its data layout, it is evident that the improvements from our multimodal data representation process are substantially more effective.
Figure 16: Range Query Evaluation Results for Different Indexes on Real Datasets.
7.3 Evaluation of KNN Query
Next, we evaluate MQRLD’s performance on KNN queries using the Yelp, Color, and Forest datasets. Similarly, under the same query workload, we optimize other indexes as much as possible. It can be observed from Fig 17 that MQRLD maintains relatively low average query times across different $K$ values (due to incomparable results beyond 12 dimensions for R*-tree and LISA, we disregard them on the Yelp and Color):
(1) Compared with non-learned indexes, MQRLD shows average query efficiency improvements of 4.2×, 3.1×, and 3.3× over M-tree, R*-tree, and SPB-tree respectively. This is because the leaf nodes of MQRLD contain highly similar data, and have a prediction model to improve the query speed.
(2) Compared with learned indexes, MQRLD improves query efficiency by an average of 2.2×, 1.9×, and 1.5× over ML, LISA, and LIMS datasets. ML uses clustering for data mapping but suffers from prediction errors and longer search times due to its global predictive model. LISA and LIMS utilize local predictive models to improve the query efficiency for "last-mile" search, but their models are complex and the mapping process is time-consuming. In contrast, MQRLD trains simple linear regression models at each leaf node, achieving more accurate and faster queries. This is enabled by MQRLD’s data-aware and query-aware data representation and high-dimensional learned index construction strategy, which keep prediction model errors within a low threshold as well as improve MQRLD’s query efficiency.
Figure 17: KNN Query Evaluation Results for Different Indexes on Real Datasets.
7.4 Evaluation of CBR
In this section, we investigate the CBR of each index on both real and synthetic datasets. Fig 18 reports the CBR for range queries across different datasets. We observe that MQRLD exhibits the lowest CBR across various datasets. The tree-based indexes, R*-tree and Qd-tree, show relatively less CBR, demonstrating that tree structure has the advantage of pruning a significant amount of irrelevant data regions. Flood uses grid indexes that also effectively filter out irrelevant data through grid partitioning, while Tsunami, which combines tree structures and grid indexes, further leverages these advantages to minimize CBR during queries. ZM and ML, however, show higher CBR across all datasets. This is due to the sequential scanning required on one-dimensional mapped values, compounded by the drawbacks of Z-order curves and prediction errors in learning models. Compared to these indexes, MQRLD combines the structural advantages of tree indexes with our index optimization strategy that ensures frequently accessed nodes are kept in positions that are more cache-friendly, allowing most queries to be completed by scanning only a few leaf nodes.
Figure 18: Cross-bucket Rate Evaluation Results of Different Indexes on Different Datasets.
Figure 19: Scalability Evaluation Results w.r.t. Dataset Size.
7.5 Scalability
Dataset Size
To show how MQRLD scales with dataset size, we create synthetic datasets at sizes of {1M,10M,50M,100M}. We train and evaluate these datasets with the same train and test workloads as the full dataset. Fig 19 shows the results of each index on different-sized datasets, where average query time for all indexes exhibits a nearly linear relationship with dataset size, with MQRLD demonstrating the best query performance overall.
In the range query experiments, for the Uniform and Skewed datasets, which lack distinct clustering characteristics, the query efficiency decreases. However, MQRLD’s data representation strategy, which involves repositioning data points in hyperspace, enhances clustering characteristics and mitigates the impact of data distribution.
In the KNN query experiments, we can observe that the query efficiency of MQRLD outperforms other indexes by 1.6× to 5.4×. Unlike other tree indexes, MQRLD’s tree depth and the number of leaf nodes do not significantly increase as the data volume increases. This allows MQRLD to effectively prune a large amount of irrelevant data during KNN queries. Additionally, the prediction models within its leaf nodes further speed up locating the lookup data points.
Figure 20: Scalability Evaluation Results w.r.t. Dimension.
Number of Dimensions
To show how MQRLD scales with dimensions, we create synthetic $d$-dimensional datasets $(d\leq 16)$ with 10 million records whose values in each dimension are distributed uniformly at random. Fig 20 shows that MQRLD continues to outperform other indexes at each dimension. The growth rates of MQRLD, LIMS, and ML are lower compared to other indexes, indicating that clustering-based data techniques effectively mitigate the curse of dimensionality phenomenon. R*-tree and LISA maintain efficient query performance in low-dimensional queries. However, their query time increases exponentially with dimensionality, and thus, their results are not recorded beyond 12 dimensions.
Both range and KNN queries on the GaussMix dataset show that the average query time for ML, LIMS, and MQRLD is lower than on other datasets, primarily due to the dataset’s inherent clustering-friendly nature. MQRLD’s data representation strategy further enhances this clustering characteristic, thereby improving its query efficiency. In range query experiments, the increase in dimensionality significantly raises the number of grids in Flood and Tsunami, which in turn escalates grid filtering time during queries. In contrast, the number of leaf nodes in MQRLD remains unaffected by the dimensionality, ensuring consistent query performance. In KNN query experiments, LIMS shows higher average query times in low-dimensional scenarios. This is due to the limited pruning information available in low dimensions, which increases query costs. Conversely, MQRLD’s pruning strategy, based on distance determination, remains stable across various dimensionality, demonstrating more robust performance.
7.6 Evaluation of Rich Hybrid Queries
In this section, we evaluate the performance of rich hybrid queries. Except for MQRLD, other indexes in our experiment do not support rich hybrid queries and can only search a single vector, so we combine them to simulate the execution of rich hybrid queries. These combinations include six different hybrid indexes: M-tree $\times$ N, SPB-tree $\times$ N, ML $\times$ N, LIMS $\times$ N, LIMS + Flood, and LIMS + Tsunami. $\times$ N means a combination of N instances of the same index (N $\in$ [2,5]). Note that these combined indexes execute queries sequentially. For example, if ML $\times$ 2 is used to perform a V.K $\bigoplus$ V.R query, we will first use one ML index to execute the V.K query, then build another ML index to perform the V.R query.
Fig 21 presents the results of the average query times for these rich hybrid queries. We can observe that MQRLD consistently achieves the lowest average query times across all experiments. In experiments for three typical rich hybrid queries (V.R $\bigoplus$ N.R, N.R $\bigoplus$ V.K, and V.R $\bigoplus$ V.K), MQRLD improves query performance by 2.8×, 1.5×, and 3.2× on average, respectively, compared to the next best-performing index, LIMS $\times$ 2. Notably, in the V.R $\times$ N experiment, while the average query time for all indexes increases as the number of columns grows, MQRLD demonstrates the smallest increase. Its query performance is, on average, 4.7× better than that of the next best-performing index, LIMS $\times$ N. This superiority is due to MQRLD’s ability to maintain a single index for multiple feature vectors, whereas other indexes require combinations to complete the query tasks. Additionally, MQRLD’s data representation strategy and index optimization algorithm are effective in rich hybrid query scenarios, optimizing data distribution and index structure based on query requirements, thus maintaining high query efficiency.
Figure 21: Rich hybrid Query Evaluation Results for Different Indexes on Real Datasets.
7.7 Discussions
In this section, to clarify some potential concerns that have not been mentioned before, we will discuss the time and space consumption associated with the proposed MQRLD and demonstrate the effectiveness of our proposed optimization strategies through ablation experiments. Furthermore, all experiments conducted to verify these concerns utilize the GuassMix dataset as the experimental data.
The Time Cost of QBS Table Construction
In this paper, the QBS table has a significant impact on the query-aware mechanism. However, since the calculation of statistics variables Recall@K and Query Accuracy in the QBS table is time-consuming, computing the statistics information corresponding to all query statements may result in significant time costs. Therefore, we sample query statements during the query process and calculate their corresponding statistics information to construct the QBS table.
The Time Cost of Feature Measurement
In feature measurement, we select the embedding model with the highest score by calculating the measurement metrics across different embedding models. However, due to the computations of SC, FID, and extrinsic measurement metric (see details in Section 5.1.2) being quite slow, performing feature measurement on the entire dataset would also result in a heavy workload. Therefore, similar to QBS table construction, we calculate the measurement metrics by sampling the data (for calculating SC and FID) and queries (for calculating extrinsic measurement metrics) to select the embedding model with the highest score.
The Time Cost of Index Construction
For a multimodal data retrieval platform, the volume of data it stores is often substantial. This raises a concern about whether constructing the MQRLD index is highly time-consuming. Experiments demonstrate that when the dataset contains 10M data, the index construction time is 1393.6s, indicating that the index construction is not excessively time-consuming.
The Space Cost of Index
In index construction, we use a linear regression model to fit the data distribution within each leaf node, resulting in each leaf node storing only a few parameters. However, with large-scale data, the number of leaf nodes may become substantial, potentially leading to an excessively large index size despite the small number of parameters per model. Fig 22(a) demonstrates that the space complexity of the index structure we constructed on datasets with varying data volumes approximates $O(logN)$. Therefore, as the amount of multimodal data stored in the platform increases, the space cost of the MQRLD index will not grow excessively.
Figure 22: The Space Cost of MQRLD Index(a), and Ablation Study(b).
Ablation Experiments
To explore the impact of different components on MQRLD performance, we evaluate the performance of range queries using MQRLD in various states. Fig 22(b) provides a detailed illustration of the average query times for different query optimized methods, in which "Full Scan" refers to scanning the entire dataset to implement range query, "Initialized_MQRLD" indicates querying with the initialized MQRLD without index optimization, and "Optimized_MQRLD" represents querying with the optimized MQRLD.
From the experimental results, we can draw the following conclusions:
(1) Comparison between "Full Scan" and "Initialized_MQRLD" shows that our data representation strategy and index structure significantly reduce query times, even in the initial state.
(2) The results for "Initialized_MQRLD" and "Optimized_MQRLD" validate the effectiveness of our query-aware mechanism, with the optimized matrix $T$ and index structure further enhancing query performance.
(3) In "Optimized_MQRLD," the sampling frequency employed is the minimum effective value for both feature enhancement and index optimization, which is 10%. As the sample frequency increases, the performance of MQRLD continues to improve.
8 Conclusion
In this study, we introduce a multimodal data retrieval platform MQRLD. This platform builds on a data lake which offers transparent data storage and integrates a multimodal open API for rich hybrid queries. In addition to these capabilities, its multimodal data representation strategy transforms data into effective feature vectors through feature embedding, measurement, and enhancement, aiding accurate data retrieval. To further improve query efficiency, high-dimensional learned indexes are built using divisive hierarchical clustering and cluster tree construction, with the query-aware mechanism optimizing the index structure. Overall, MQRLD not only supports transparent data storage and rich hybrid queries, but also ensures effective and efficient retrieval.
Our study provides a broad scope for future work in multimodal data management. Additionally, the optimization strategies we propose for the multimodal data retrieval platform also have versatile applications. Experiments have demonstrated that our proposed multimodal data representation strategy and high-dimensional learned index effectively improve retrieval accuracy and efficiency on data lakes. Despite the limitations of existing platforms in multimodal data retrieval, their inherent features offer potential in specific scenarios, as discussed in Section 1. Therefore, we consider applying our multimodal data representation strategy and high-dimensional learned index to other platforms in the future to enhance their multimodal data retrieval performance.
[…]
Acknowledgements
The work is funded by the National Natural Science Foundation of China under Grant 62076027.